# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ 4. Exponential Jordan type
(4.3.3) $\mathcal{U}_{\mathfrak{g},r|s}\ \equiv\ \mathcal{E}_{r|s}\times pr_{2},\quad\mathcal{U}_{\mathfrak{g},r}\ \equiv\ \mathcal{E}_{\mathfrak{g},r}\times pr_{2},$
the “scalar extensions” of $\mathcal{E}_{r|s},\ \mathcal{E}_{(r)}$ along $k\to k[\mathcal{C}_{r}(\mathfrak{g})]$.
Remark 4.4.
Observe that for any scheme-theoretic point $\underline{B}=(B_{0},\ldots,B_{r-1})$, $(\mathcal{E}_{r|s})_{*}(u_{r-1})$ determines
(4.4.1) $t^{p^{r-1}}\in\mathbb{G}_{a(r)}(K[t]/t^{p^{r}})\ \mapsto\ t^{p^{r-s-1}}\in\mathbb{G}_{a(r)}(K[t]/t^{p^{r}})\ \mapsto(\mathcal{E}_{(r),B_{s}})_{*}(t^{r-s-1})\in\mathbb{G}_{(r)}(K[t]/t^{p^{r}})$
where $K=k(\underline{B})$ and we identify $t^{p^{r-1}}\in\mathbb{G}_{a(r)}(k[t]/t^{p^{r}})=k[t]/t^{p^{r}}$ with $u_{r-1}\in(k[t]/t^{p^{r}})^{*}$. (See [22, Eqn 1.3.1] with $S=k[t]/t^{p^{r}}$.)
Thus, $(\mathcal{E}_{\mathfrak{g},r})_{*}(u_{r-1})$ determines
$t^{p^{r-1}}\quad\mapsto\quad\prod_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(t^{r-s-1})\ \in\ \mathbb{G}_{(r)}(K[t]/t^{p^{r}}).$
The following proposition “specializes” the structures of Definition 2.8 for an affine group scheme $\mathbb{G}$ satisfying the condition that it is equipped with $\mathcal{E}{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}(\mathfrak{g})\to\mathbb{G}_{(r)}$ as in Definition 4.2.
@@
    $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}:\mathbb{G}_{a(r),k[\mathcal{C}_{r}(\mathfrak{g})]}\quad\to\quad\mathbb{G}_{(r),k[\mathcal{C}_{r}(\mathfrak{g})]}$
@@
    $\Phi_{\mathfrak{g},r|s},\ \Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\quad\to\quad V_{r}(\mathbb{G})$
    with the property that their pull-backs of the universal 1-parameter group  $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\to\ \mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$ equal  $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}$.
@@
    $\Phi_{\mathfrak{g},r}^{*}(\mathcal{U}_{\mathbb{G},r})\ =\ \mathcal{U}_{\mathfrak{g},r}$, where $\mathcal{U}_{\mathbb{G},r}$ is given in Definition 2.9.
  4. (4)
    Each $\Phi_{\mathfrak{g},r|s}$ commutes with the monoid actions of $\mathbb{A}^{1}$ on $\mathcal{C}_{r}(\mathfrak{g})$ and $V_{r}(\mathbb{G})$ in the sense that the follow squares commute:
[…]
@@
Over any point $\underline{B}$ of $\mathcal{C}_{r}(\mathfrak{g})$, $\mathcal{U}_{\mathfrak{g},r|s}and\ \mathcal{U}_{\mathfrak{g},r}$ are morphisms of group schemes over $\underline{B}$; this implies assertion (1). Assertions (2) and (3) follow from the universal property of $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\quad\to\quad\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$.
@@
In [22], the authors introduce the condition that an embedding $\mathbb{G}\hookrightarrow GL_{N}$ of linear algebraic groups is of exponential type; they observe that many classical groups satisfy this condition, including many classical simple groups, their parabolic subgroups, and unipotent radicals of these parabolic subgroups. Subsequent generalizations of this condition on $\mathbb{G}$ occur in [10] and [11]. A more definitive account of such linear algebraic groups can be found in P. Sobaje’s article [21]. The following definition is general enough to include examples of those papers.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. Assume that $\mathbb{G}$ is equipped with a morphism
satisfying the conditions of Definition 4.2. We say that $\mathbb{G}$ equipped with $\mathcal{E}_{(r)}$ is of exponential type of height $r$ if the morphism $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\ \to\ V_{r}(\mathbb{G})$ of Theorem 4.5(1) induces a bijection on geometric points.
In [9, Defn 1.6], the notions of a “continuous algebraic map” and a “bicontinuous morphism” were introduced in order to study rational maps between schemes which induce well defined maps on geometric points. In that work, the topologies of interest were either the ètale topology or the analytic topology (in the case that the ground field was the complex numbers).
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be affine group scheme of exponential type of height $r$. The bicontinuous morphism
(4.7.1) $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\quad\to\quad V_{r}(\mathbb{G}).$
@@
The assumption of exponential type of height $r$ implies that $\Phi$ is a bijection on scheme-theoretic points. Since $\Phi_{\mathfrak{g},r}$ is a morphism, it is continuous. Since $\mathbb{P}\Phi_{\mathfrak{g},r}:\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\to\mathbb{P}V_{r}(\mathbb{G})$ is proper, it is a closed map, thus inducing a bijection on closed subsets, thus a homeomorphism. This implies that the pull-back/restriction $\mathcal{C}_{r}(\mathfrak{g})-\{0\}\to V_{r}(\mathbb{G})-\{0\}$ is a homeomorphism. Since the cone point 00 is closed in both $\mathcal{C}_{r}(\mathfrak{g})$ and $V_{r}(\mathbb{G})$, this implies that $\Phi_{\mathfrak{g},r}$ is also a homeomorphism. ∎
As in (2.9.1) of Theorem 2.11, we consider $p$-nilpotent operators associated to the 1-parameter subgroups  $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}$ of Proposition 4.5(2).
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be affine group scheme of exponential type of height $r$. We define the $p$-nilpotent operators
(4.8.1) $\Theta_{\mathfrak{g},r|s}\ \equiv\ (\mathcal{U}_{\mathfrak{g},r|s})_{*}(1\otimes u_{r-1}),\ \Theta_{\mathfrak{g},r}\ \equiv\ (\mathcal{U}_{\mathfrak{g},r})_{*}(1\otimes u_{r-1})\ \in\ k[\mathcal{C}_{r}(\mathfrak{g})]\otimes\mathbb{G}_{(r)},$
and denote by $\Theta_{\mathfrak{g},r|s,\underline{B}},\ \Theta_{\mathfrak{g},r,\underline{B}}$ their specializations at a scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$.
@@
    $\Theta_{\mathfrak{g},r|s,\underline{B}}\ =\ (\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1}),\ \Theta_{\mathfrak{g},r,\underline{B}}\ =\prod_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}\quad\in\ k(\underline{B})\otimes\mathbb{G}_{(r)}$
@@
    Both $\Theta_{\mathfrak{g},r|s}^{exp}$ and $\Theta_{\mathfrak{g},r}^{exp}$ are homogeneous of degree $p^{r-1}$ with respect to the grading on $k[\mathcal{C}_{r}(\mathfrak{g})]$.
@@
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\Theta_{\mathfrak{g},r|s}$ and $\Theta_{\mathfrak{g},r}$ determines $p$-nilpotent, $k[\mathcal{C}_{r}(\mathfrak{g})]$-linear operators
    $\Theta_{\mathfrak{g},r|s,M},\ \Theta_{\mathfrak{g},r,M}:k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M\ \to\ k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M.$
@@
    The base changes along a scheme theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$ of the operators $\Theta_{\mathfrak{g},r,|sM},\ \Theta_{\mathfrak{g},r,M}$ are given by the actions of $\Theta_{\mathfrak{g},r,|s\underline{B}},\ \Theta_{\mathfrak{g},r,\underline{B}}$ on $k(\underline{B})\otimes M$.

### v2

@@ 4. Exponential Jordan type
(4.3.3) $\mathcal{U}_{\mathfrak{g},r|s}\equiv\mathcal{E}_{r|s}\times pr_{\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))},\ \mathcal{U}_{\mathfrak{g},r}\equiv\mathcal{E}_{\mathfrak{g},r}\times pr_{\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}:\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathbb{G}_{(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})),$
given by maps on coordinate agebras as the the $k[\mathcal{C}_{r}(\mathfrak{g})]$-linear extensions of $\mathcal{E}_{r|s}^{*},\ \mathcal{E}_{\mathfrak{g},r}^{*}$ along $k\to k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$.
The following proposition identifies the specializations of the maps $\mathcal{U}_{\mathfrak{g},r|s}$ and $\mathcal{U}_{\mathfrak{g},r}$ at a geometric point of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$. In particular, the specialization of $\mathcal{U}_{\mathfrak{g},r}$ at a geometric point $\underline{B}=(B_{0},\ldots,B_{r-1})$ of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is expressed as a product of Frobenius twists of exponentiations of the $B_{i}$’s.
Proposition 4.4.
Retain the hypotheses and notation of Definition 4.1. For any geometric point $\underline{B}=(B_{0},\ldots,B_{r-1})$ of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$, the specialization of $\mathcal{U}_{\mathfrak{g},r|s}$ at $\underline{B}$ is denoted by $\mathcal{U}_{\mathfrak{g},r|s,B_{s}}$. Then the specialization of $\mathcal{U}_{\mathfrak{g},r}$ at $\underline{B}$ is given by
(4.4.1) $\mathcal{U}_{\mathfrak{g},r,\underline{B}}\ =\ \prod_{s=0}^{r-1}\mathcal{U}_{\mathfrak{g},r|s,B_{s}}\circ F^{s}:\mathbb{G}_{a(r),k(\underline{B})}\ \to\mathbb{G}_{a(r),k(\underline{B})}\ \to\ \mathbb{G}_{(r),k(\underline{B})}.$
In particular,
(4.4.2) $(\mathcal{U}_{\mathfrak{g},r,\underline{B}})_{*}:k(\underline{B})\otimes k\mathbb{G}_{a(r)}\to k(\underline{B})\otimes k\mathbb{G}_{(r)},\quad 1\otimes u_{r-1}\ \mapsto\ \prod_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(u_{r-s-1}).$
Proof.
Equations (4.4.1) and (4.4.1) follow from equations (4.3.1) and (4.3.2) together with the definitions of $\mathcal{U}_{\mathfrak{g},r|s,\underline{B}}$ and $\mathcal{U}_{\mathfrak{g},r,\underline{B}}$ as $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$-linear extensions of $\mathcal{E}_{r|s},\ \mathcal{E}_{\mathfrak{g},r}$.
Equation (4.4.2) follows immediately from (4.4.1). ∎
We proceed to investigate the close relationship between $\mathcal{U}_{\mathbb{G},r}$ and $\mathcal{U}_{\mathfrak{g},r}$.
@@
    $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}:\mathbb{G}_{a(r),k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]}\quad\to\quad\mathbb{G}_{(r),k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]}$
@@
    $\Phi_{\mathfrak{g},r|s},\ \Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad V_{r}(\mathbb{G})$
    with the property that their pull-backs of the universal 1-parameter group $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\to\ \mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$ equal the 1-parameter subgroups $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}$.
@@
    The morphisms $\Phi_{\mathfrak{g},r|s},\Phi_{\mathfrak{g},r}$ commute with the monoid actions of $\mathbb{A}^{1}$ on $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ and $V_{r}(\mathbb{G})$ in the sense that each of the following squares commute:
    (4.5.1) ${\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\times\mathbb{A}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\mathfrak{g},r|s}\times 1}$${\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{r,s}}$${V_{r}(\mathbb{G})\times\mathbb{A}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V_{r}(\mathbb{G}).}$
@@
As can be seen using Proposition 4.4 and the condition of Definition 4.1(2), $\mathcal{U}_{\mathfrak{g},r|s,\underline{B}}$ and $\mathcal{U}_{\mathfrak{g},r,\underline{B}}$ are morphisms of group schemes over $\underline{B}$. This implies assertion (1). Assertion 2 follows from the universal property of $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\quad\to\quad\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$.
@@
In [26], Suslin, Bendel, and the author introduced the condition that an embedding $\mathbb{G}\hookrightarrow GL_{N}$ of linear algebraic groups be of exponential type; they observe that various algebraic groups satisfy this condition, including many classical simple groups, their parabolic subgroups, and unipotent radicals of these parabolic subgroups. Subsequent generalizations of this condition on $\mathbb{G}$ occur in [12] and [13]. A more definitive account of such linear algebraic groups can be found in P. Sobaje’s article [25]. The following definition is a natural generalization of those earlier considerations of groups of exponential type. The motivating special case is the exponential function of $GL_{N}$ applied to a $p$-nilpotent element of $\mathfrak{g}l_{N}$.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ equipped with an exponentiation $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\ \to\ \mathbb{G}_{(r)}$ of height $\leq r$. We say that $\mathbb{G}$ equipped with $\mathcal{E}_{(r)}$ is of exponential type of height $r$ if the morphism $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\ \to\ V_{r}(\mathbb{G})$ of Theorem 4.5(1) induces a bijection on geometric points.
In [11, Defn 1.6], the notions of a “continuous algebraic map” and a “bicontinuous morphism” were introduced in order to study rational maps between schemes which induce well defined maps on geometric points. In that work, the topologies of interest were either the ètale topology or the analytic topology (in the case that the ground field was the complex numbers). We are now working with varieties over $k$, a field of characteristic $p>0$.
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. The bicontinuous morphism
(4.7.1) $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad V_{r}(\mathbb{G})$
@@
The assumption of exponential type of height $r$ implies that $\Phi_{\mathfrak{g},r}$ is a bijection on scheme-theoretic points. Since $\Phi_{\mathfrak{g},r}$ is a morphism, it is continuous. Since $\mathbb{P}\Phi_{\mathfrak{g},r}:\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathbb{P}V_{r}(\mathbb{G})$ is proper, it is a closed map, thus induces a bijection on closed subsets, thus a homeomorphism. This implies that the pull-back/restriction $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))-\{0\}\to V_{r}(\mathbb{G})-\{0\}$ is a homeomorphism. Since the cone point 0 is closed in both $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ and $V_{r}(\mathbb{G})$, this implies that $\Phi_{\mathfrak{g},r}$ is also a homeomorphism. ∎
We adapt the construction of $\Theta_{\mathbb{G},r}$ in Definition 2.13 and its use in Theorem 2.15 by replacing $\mathcal{U}_{\mathbb{G},r}$ by $\mathcal{U}_{\mathfrak{g},r}$.
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. We define the $p$-nilpotent operators
(4.8.1) $\Theta_{\mathfrak{g},r|s}\ \equiv\ (\mathcal{U}_{\mathfrak{g},r|s})_{*}(1\otimes u_{r-1}),\ \Theta_{\mathfrak{g},r}\ \equiv\ (\mathcal{U}_{\mathfrak{g},r})_{*}(1\otimes u_{r-1})\ \in\ k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})))]\otimes k\mathbb{G}_{(r)},$
and denote by $\Theta_{\mathfrak{g},r|s,\underline{B}},\ \Theta_{\mathfrak{g},r,\underline{B}}$ their specializations at a scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$.
@@
    $\Theta_{\mathfrak{g},r|s,\underline{B}}\ =\ (\mathcal{E}_{r|s,B_{s}})_{*}(u_{r-s-1}),\ \Theta_{\mathfrak{g},r,\underline{B}}\ =\prod_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}\quad\in\ k(\underline{B})\otimes k\mathbb{G}_{(r)},$ elements of the group algebra of $\mathbb{G}_{(r),k(\underline{B})}$.
@@
    Both $\Theta_{\mathfrak{g},r|s}$ and $\Theta_{\mathfrak{g},r}$ are homogeneous of degree $p^{r-1}$ with respect to the grading on $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$.
@@
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\Theta_{\mathfrak{g},r|s}$ and with $\Theta_{\mathfrak{g},r}$ determines $p$-nilpotent, $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$-linear operators
    $\Theta_{\mathfrak{g},r|s,M},\ \Theta_{\mathfrak{g},r,M}:k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M\ \to\ k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M.$
@@
    The base changes along a scheme theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of the operators $\Theta_{\mathfrak{g},r|s,M},\ \Theta_{\mathfrak{g},r,M}$ are given by the actions of $\Theta_{\mathfrak{g},r|s,\underline{B}},\ \Theta_{\mathfrak{g},r,\underline{B}}$ on $k(\underline{B})\otimes M$.
@@
Observe that the elements $\Theta_{\mathfrak{g},r|s},\ \Theta_{\mathfrak{g},r}$ are $p$-nilpotent because $1\otimes u_{r-1}\in\mathbb{G}_{a(r),\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]}$ is $p$-nilpotent. Assertion (1) follows from Proposition 4.5(2), the definition of $\Theta_{\mathbb{G},r}$ in Definition 2.13, and the definition of $\Theta_{\mathfrak{g},r}$ given above. The first part of assertion (2) identifying $\Theta_{\mathfrak{g},r|s,\underline{B}}$ follows from Proposition 4.4, so that the second part is verified using the fact that specialization along $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\to k(\underline{B})$, $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes\mathbb{G}_{(r)}\to k(\underline{B})\otimes\mathbb{G}_{(r)}$, commutes with the product map of $\mathbb{G}_{(r)}$.